OpenStudy (anonymous):

Part 1: Explain how you calculate the restricted value(s) of x for the equation and what those value(s) represent. Part 2: Explain which method (cross-multiplication or LCD) you would use to solve the equation and why. Part 3: Using complete sentences, explain the steps to solve the equation using the method you chose in step 2. Provide the solution to the equation.

OpenStudy (anonymous):

\[\frac{ 3}{ x+2 } + \frac{ 1 }{ 2x } = \frac{ 4 }{ x+2 }\] thats the equation

OpenStudy (anonymous):

did you pray?

OpenStudy (anonymous):

what?